OpenStudy (anonymous):

is (-6,-4) a solution to the equation 6x+3y=19?

OpenStudy (karatechopper):

plug it in! what do u think?

OpenStudy (karatechopper):

nobody answer let him try by himself

OpenStudy (amistre64):

lol .... anybody can answer ...

OpenStudy (karatechopper):

yes i know but i want him to try himself

OpenStudy (amistre64):

when (x=-6, y=-4) do we get a true statement?

OpenStudy (anonymous):

can ya please help me with that problem please for the linear equation x-2y=10 which of the following has 10 a.slope b. the x-intercept c. the y-intercept d. the orgin

OpenStudy (anonymous):

true statment

OpenStudy (karatechopper):

hey amistre i umm am thinking something else what u think?

OpenStudy (amistre64):

please dont spam other peoples posts with your own questions; it comes across as a bit rude.

OpenStudy (amistre64):

6x+3y=19 -6 -4 -36-12=19 ..... im thinking its something else too

OpenStudy (karatechopper):

thank u!